A. If you do 12 J of work to push 0.001 C of charge from point A to point B in an electric field, what is the potential difference between points A and B?

 

B. What will be the increase in kinetic energy of an electron if it has been accelerated through a potential difference of 20 million volts? (Assume that e = 1.6 X 10-19 C)


1
Expert's answer
2022-03-27T18:21:03-0400

A.

"V=\frac{W}{q}=\frac{12\:\rm J}{0.001\:\rm C}=12000\:\rm V"

B.

"\Delta E_k=qV\\
=1.6*10^{-19}*20*10^6=3.2*10^{-12}\:\rm J"